OpenStudy (anonymous):

What is the equation of the line passing through (2, -3) and parallel to the line y = -6x - 1 in standard form?

OpenStudy (anonymous):

a) 6x + y = 9 b) 6x - y = -9 c) x + 6y = 9 d) x - 6y = -9

OpenStudy (shamim):

the equation of line will b\[y-y _{1}=m(x-x _{1})\]

OpenStudy (shamim):

u hv slope of ur given line is\[m=-6\]

OpenStudy (shamim):

so any line which is parallel to ur given line hv slope\[m=-6\]

OpenStudy (anonymous):

so its y-(-3)=-6(x-2)

OpenStudy (shamim):

so the equation of ur required line will b\[y-(-3)=-6(x-2)\]

OpenStudy (shamim):

ya

OpenStudy (anonymous):

i got b. 6x-y=-9

OpenStudy (anonymous):

am i correct?

OpenStudy (anonymous):

formula => \[(y-y1)=m(x-x1)\] use this formula where m=-6 and (x1 , y1)=(2,-3) u will get the equation

OpenStudy (shamim):

may i know the name of ur country @danny562

OpenStudy (anonymous):

ans is option a ....

OpenStudy (anonymous):

wait is the answer a or b?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

6x+y=9

OpenStudy (anonymous):

s.... solve by substituting ...

OpenStudy (anonymous):

noo ....y+6x=9

OpenStudy (anonymous):

\[y-(-3)=-6(x-2) \] \[y+3-6x+12\] \[y+6x=12-3\] \[y+6x=9\]

OpenStudy (anonymous):

in options its 6x+y=9 same as y+6x=9